a series circuit with R=10 ohms,L=100mH and C=100microF is supplied with 200 V/50Hz. Find the impedance,current,power,power factor.


1
Expert's answer
2021-04-23T07:25:40-0400

Impedance

"X_L=2\pi fL=2\pi \times 50 \times 100 \times 10^{-3}=20\pi"

"X_C=\frac{1}{2\pi fL}=\frac{1}{2\pi \times 50 \times 100 \times 10^{-3}}=\frac{1}{20\pi}"

"Z=\sqrt{10^2+(20\pi- \frac{1}{20\pi}})^2=63.6 \Omega"

Current

"I=\frac{V}{Z}= \frac{200}{63.6}=3.145 A"

Power

"P=IV=3.145 \times 200=628.93 W"

Power factor

"cos \phi =\frac{R}{Z}=\frac{10}{63.6}=0.157"